A normal distribution has mean 63 and standard deviation 5. Find the data value if the z-score is .22
Answer by jim_thompson5910(35256)

z = (x - mu)/sigma

0.22 = (x - 63)/5

5*0.22 = x - 63

1.1 = x - 63

1.1 + 63 = x

64.1 = x

x = 64.1

The data value is 64.1